# Chinese Academy of Sciences, Beijing 100190, China

### v1

@@ 3.1 QKEs in the early universe
In the early universe, both the Standard Model (SM) thermal bath and the neutrino system can be approximated as homogeneous and isotropic. Under these conditions, the QKEs for neutrinos can be significantly simplified. First, the drift and force terms on the left-hand side are eliminated, and a Hubble term is introduced to account for the expansion of the Universe. Second, the density matrix for neutrinos is considered to depend solely on the absolute value of the momentum. Finally, the coherence matrices between left-handed and right-handed neutrinos, namely $\varrho_{LR,RL}$ and $\bar{\varrho}_{LR,RL}$ for Dirac neutrinos, and $\phi$ for Majorana neutrinos, are reduced to zero, leading to corresponding collision terms also being zero. Consequently, the non-zero density matrices are $\varrho_{LL,RR}$ and $\bar{\varrho}_{LL,RR}$ for Dirac neutrinos, and $\varrho$ and $\bar{\varrho}$ for Majorana neutrinos.
For active neutrinos, the QKEs for $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ in the case of Dirac neutrinos, are identical with those for $\varrho$ and $\bar{\varrho}$ for Majorana neutrinos. Therefore, in the subsequent analysis, we will not distinguish between Dirac and Majorana neutrinos and will use $\varrho$ and $\bar{\varrho}$ to represent the density matrices for neutrinos and antineutrinos, respectively. In the flavor basis, the QKEs can be written as:
@@
Here, $\mathbb{M}$ represents the neutrino mass matrix in the mass basis, and $U$ is the PMNS transformation matrix that relates the flavor eigenstates to the mass eigenstates. Using the standard parametrization of the PMNS matrix Giunti:2007ry ; Gariazzo:2019gyi , we have:
@@
where $c_{ij}=\cos{\theta_{ij}}$, $s_{ij}=\sin{\theta_{ij}}$, and $\theta_{ij}$ are the mixing angles. We neglect the potential CP-violating phase in this work.
The remaining terms in the Hamiltonian account for the matter potentials induced by coherent scatterings between neutrinos and other plasma particles. Here, $m_{W}$ and $m_{Z}$ denote the masses of the $W$ and $Z$ bosons, respectively, and $G_{F}$ is the Fermi constant. The second term in the Hamiltonian reflects the net number densities for charged leptons ($\mathbb{N}_{l}\equiv\mathrm{diag}(n_{e}-n_{\bar{e}},n_{\mu}-n_{\bar{\mu}},0)$) and neutrinos ($\mathbb{N}_{\nu}\equiv n_{\nu}-n_{\bar{\nu}}$). The third term accounts for the energy densities of charged leptons and neutrinos ($\mathbb{E}_{l}\equiv\mathrm{diag}(\rho_{e}+\rho_{\bar{e}},\rho_{\mu}+\rho_{\bar{\mu}},0)$, $\mathbb{P}_{l}\equiv\mathrm{diag}(P_{e}+P_{\bar{e}},P_{\mu}+P_{\bar{\mu}},0)$, $\mathbb{E}_{\nu}\equiv\rho_{\nu}+\rho_{\bar{\nu}}$). In practice, the net number densities for charged leptons are often negligible due to the small baryon asymmetry. However, the net number densities for neutrinos can lead to significant effects, especially if primordial neutrino asymmetries are present. This can result in pronounced neutrino flavor oscillations, including the well-known "synchronized oscillation" when this term dominates the Hamiltonian Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa ; Johns:2016enc ; Froustey:2021azz ; Froustey:2024mgf .
Finally, the collision terms $\mathcal{I}$ and $\bar{\mathcal{I}}$ account for contributions from neutrino-electron/positron scattering ($\nu e\leftrightarrow\nu e$), annihilation ($\nu\bar{\nu}\leftrightarrow e^{+}e^{-}$), and neutrino-neutrino self-interactions ($\nu\nu\leftrightarrow\nu\nu$, $\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$). These terms can be expressed as
@@
with a similar decomposition for the collision term of antineutrinos $\bar{\mathcal{I}}$. The collision terms are generally functions of the neutrino density matrices, and their explicit forms are detailed in Appendix A. Contributions from other interaction processes, such as $\mu^{\pm}$ scattering and annihilation, are neglected due to their minimal impact at the temperature of neutrino decoupling.
In practice, comoving coordinates and corresponding comoving variables are employed to eliminate the Hubble expansion term in the QKEs. The comoving temperature is defined as $T_{\rm CM}\propto a^{-1}$ Grohs:2015tfy , with $T_{\rm CM}=T_{\nu}=T_{\gamma}$ at the initial time before neutrino decoupling when all species are tightly coupled. Using the comoving temperature, the reduced scale factor, the comoving momentum, and the dimensionless photon temperature are expressed as Mangano:2005cc ; Froustey:2020mcq ; Bennett:2020zkv
@@
By definition, the dimensionless photon temperature $z$ equals 1 at high temperatures and increases beyond 1 later due to the annihilation of electrons and positrons into photons. For the convenience of the QKEs, we also define the dimensionless energy density and pressure as $\tilde{\rho}\equiv(x/m_{e})^{4}\rho$ and $\tilde{P}\equiv(x/m_{e})^{4}P$, respectively.
The comoving QKEs for neutrinos and antineutrinos then take the form as:
@@
$\frac{{\rm d}\bar{\varrho}(x,y)}{{\rm d}x}$ $=$ $\sqrt{\frac{3m^{2}_{\rm Pl}}{8\pi\rho}}\left\{-i\frac{x^{2}}{m_{e}^{3}}\left[\tilde{H}_{\bar{\nu}},\varrho\right]+\frac{m_{e}^{3}}{x^{4}}\tilde{\bar{\mathcal{I}}}(\bar{\varrho},\varrho)\right\}\,,$
with the comoving Hamiltonian terms
@@
Additionally, $\tilde{\mathcal{I}}$ and $\tilde{\bar{\mathcal{I}}}$ represent the comoving collision terms for neutrinos and antineutrinos, respectively. The computation of these collision terms is the most complex and time-consuming aspect of solving the QKEs. Fortunately, the nine-dimensional collision integrals can be simplified to two-dimensional integrals by exploiting the homogeneous and isotropic conditions and the specific forms of the scattering amplitudes Hannestad:1995rs ; Dolgov:1997mb ; Grohs:2015tfy ; Blaschke:2016xxt . The explicit forms of the reduced collision terms are provided in Appendix A.
Furthermore, as noted in the discussion following the QKEs (Eq. (53)), the chemical potential of electrons and the $\mathbb{N}_{l}$ term in the Hamiltonian are negligible due to ue to the very low baryon-to-photon ratio $\eta$ and the efficient electromagnetic interactions. Consequently, the Hamiltonian is predominantly influenced by the net number densities of neutrinos when primordial lepton asymmetries are sufficiently large and vary among different flavors. This results in synchronized flavor oscillations of neutrinos and antineutrinos Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa . Consequently, neutrino asymmetries tend to equilibrate among different flavors prior to the epoch of BBN Froustey:2021azz ; Froustey:2024mgf . Since this work focuses on the implications of neutrino asymmetries for BBN, CMB, and LSS, we will concentrate on flavor-equilibrated primordial asymmetries, i.e., the initial primordial asymmetries are set to $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{n}$, where $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T_{\rm CM}$ is the degeneracy parameter for neutrinos and antineutrinos of flavor $\alpha$. By adopting this initial condition, we bypass the intricate mapping between arbitrary primordial neutrino asymmetries and those after flavor equilibration, leaving the investigation of this mapping for future work. Recent study on such mapping and the implications of distinct neutrino asymmetries among flavors for BBN can be found in  Froustey:2024mgf .
On the other hand, it is well-known that the annihilation of electrons and positrons shortly after neutrino decoupling partially heats neutrinos, leading to a slightly increased $N_{\rm eff}$. To accurately determine $N_{\rm eff}$, finite-temperature QED (FTQED) corrections must be considered Fornengo:1997wa ; Mangano:2001iu ; Bennett:2019ewm . In this work, we address three aspects of FTQED corrections: (1) Corrections to the total energy pressure densities, which primarily affect the Hubble rate; (2) Corrections to the electron mass, which primarily influence the collision terms in the QKEs; (3) Corrections to the continuity equation of the Universe, typically expressed as a differential equation for the dimensionless photon temperature $z$,
@@
where $r=x/z$, $\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}$, and $\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}$ are the diagonal components of the QKEs for neutrinos and antineutrinos. The explicit forms of the remaining $J_{i}(r)$ and $G_{i}(r)$ functions appearing in the equation, as well as other FTQED corrections, can be found in Appendix B. In practice, this equation should be solved in conjunction with the QKEs, and we consider corrections up to order $\mathcal{O}(e^{3})$ while neglecting logarithmic terms in the FTQED corrections as their effects are minimal Bennett:2019ewm ; Bennett:2020zkv ; further details are provided in Appendix B. Other FTQED corrections, such as those to the interaction vertices, are omitted in this work.
Furthermore, we restrict our analysis to the normal mass hierarchy of neutrinos and neglect any possible CP-violating phase in the transformation matrix between the neutrino flavor and mass bases. Based on the most recent values from the Particle Data Group pdg , we adopt the following neutrino mass differences and mixing angles: $\Delta m_{21}^{2}=7.53\times 10^{-5}\,\text{eV}^{2}$, $\Delta m_{31}^{2}=2.53\times 10^{-3}\,\text{eV}^{2}$, $\sin^{2}\theta_{12}=0.307$, $\sin^{2}\theta_{23}=0.545$, and $\sin^{2}\theta_{13}=0.0218$. The results for neutrino decoupling and BBN are independent of the minimal neutrino mass, provided $m_{1}\ll 0.1\,\text{MeV}$. On the other hand, the resulting CMB anisotropies and LSS of the Universe are sensitive to the choice of $m_{1}$, which will be discussed in more detail in Sections 5 and 6.
@@ 3.2 Numerical methods
To solve the QKEs for neutrinos and antineutrinos numerically, we have extended the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , initially designed for solving the QKEs in the Standard Model case, to accommodate primordial neutrino asymmetries. This subsection outlines the key numerical methods employed in the code.
For the momentum grid, the comoving momentum $y$ is discretized on a linear grid with $y_{\rm min}=0.01$, $y_{\rm max}=20$, and a total of $N_{y}=40$ points. The integrals are computed using the Newton-Cotes (NC) formula of order 1. The initial conditions are specified with an initial reduced scale factor $x=0.005$, corresponding to a comoving temperature $T_{\rm CM,in}\simeq 102.2\text{ MeV}$. The initial common temperature for all species is computed by conserving total entropy from an even higher temperature, yielding $z_{\rm in}=2.81\times 10^{-6}$. The initial density matrices are given by
$\varrho(x_{\rm in},y)={\rm diag}\{f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y)\},\quad\bar{\varrho}(x_{\rm in},y)={\rm diag}\{f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y)\}.$ (60)
@@
$f_{\nu}^{({\rm in})}(y)\equiv\frac{1}{e^{(y-\xi_{\nu})/z_{\rm in}}+1}\quad\text{and}\quad f_{\bar{\nu}}^{({\rm in})}(y)\equiv\frac{1}{e^{(y+\xi_{\nu})/z_{\rm in}}+1}\,.$ (61)
Finally, the QKEs are solved together with the energy conservation equation for $z$ using the DLSODA method from the ODEPACK hindmarsh1982odepack Fortran library. The absolute and relative error tolerances are set to $10^{-7}$, ensuring that the numerical errors of the results remain below $10^{-5}$.
@@ 3.3 Numerical results of neutrino decoupling
We present the numerical results of neutrino decoupling in this subsection. First, we consider the final energy density of neutrinos, which is typically parameterized by the effective number of relativistic species $N_{\rm eff}$, defined as
$N_{\rm eff}\equiv\frac{87}{8}\left(\frac{11}{4}\right)^{4/3}\frac{\rho_{\nu}}{\rho_{\gamma}}\,,\quad\rho_{\nu}=\sum_{\alpha}\int\frac{d^{3}p}{(2\pi)^{3}}\,p\,(\varrho_{\nu_{\alpha}}+\bar{\varrho}_{\nu_{\alpha}})$ (62)
where $\rho_{\nu}$ represents the total neutrino energy density obtained by solving the QKEs and $\rho_{\gamma}$ is the photon energy density. For the SM in the instantaneous decoupling limit, $N_{\rm eff}$ equals the number of neutrino flavors. However, since the temperature of neutrino decoupling is close to that of electron-positron annihilation, neutrinos are also heated through weak interactions with electrons and positrons, resulting in a slightly larger $N_{\rm eff}$. The most accurate calculation for the SM $N_{\rm eff}$ to date, which includes neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$ order, and neutrino-electron and neutrino-neutrino collision terms, yields $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
Figure 2: Left panel: Results of the neutrino non-instantaneous decoupling correction $\delta N_{\rm eff}^{\rm dec}=N_{\rm eff}-N_{\rm eff}^{\rm inst}$ as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The solid and dashed curves represent the results of the numerical solution and the fitting function Eq. (63), respectively. Right panel: Relative differences in the number densities of electron neutrinos and antineutrinos compared to the SM predictions in the instantaneous decoupling limit, $n_{\nu_{e},\bar{\nu}_{e}}\simeq 56\,\text{cm}^{-3}$.
In the presence of non-zero primordial neutrino asymmetries, $N_{\rm eff}$ receives additional contributions from the neutrino chemical potentials. In the instantaneous decoupling limit, the comoving distribution of neutrinos and antineutrinos of flavor $\alpha$ can be approximated as $f_{\nu_{\alpha},\bar{\nu}_{\alpha}}=\frac{1}{e^{y\,\pm\,\xi_{\nu_{\alpha}}}+1}$. Therefore, the total $N_{\rm eff}$ can be estimated by
$N_{\rm eff}^{\rm inst}\simeq 3+\sum\limits_{\alpha}\left(\frac{30}{7\pi^{2}}\xi_{\nu_{\alpha}}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu_{\alpha}}^{4}\right)=3+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right),$ (63)
where we have used the condition $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$ in the second equality.
@@
Next we present the resulting density matrices for neutrinos and antineutrinos after decoupling. The density matrices as a function of the comoving momentum $y\equiv p/T_{\rm CM}$ in the flavor basis and the mass basis for several choices of $\xi_{\nu}$ are shown in the left and right panels of Fig. 3 respectively, normalized to the Fermi-Dirac distribution with $\xi_{\nu}=0$. The density matrices of neutrinos and antineutrinos with $\xi_{\nu}=0$ coincide with each other and are represented by the gray curves. Our results also reproduce the previous density matrices obtained in the literature. For each choice of In the case of non-zero $\xi_{\nu}$, the density matrices of neutrinos and antineutrinos no longer coincide and are represented by red and green curves, respectively. For each case we use the solid, dashed and dash-dotted curve to indicate the three flavor eigenstates ($\nu_{e}$, $\nu_{\mu}$ and $\nu_{\tau}$) and the three mass eigenstates ($\nu_{1}$, $\nu_{2}$ and $\nu_{3}$), respectively. For each choice of the $\xi_{\nu}$, the corresponding thermal equilibrium distributions for neutrinos and antineutrinos $f_{\rm FD}^{\nu,\bar{\nu}}=1/(e^{y\,\pm\,\xi_{\nu}}+1)$ are also plotted by the dotted curves. Clearly, the density distributions of neutrinos and antineutrinos for the given $\xi_{\nu}$ coincide with the corresponding Fermi-Dirac distributions when $y$ is small, while the spectra are distorted at large $y$. This is due to the fact that the weak interaction rate between neutrinos and electrons/positrons is proportional to the neutrino momentum, so that the neutrino with smaller momentum decouples earlier, preserving the thermal Fermi-Dirac distribution, while the neutrino with larger momentum decouples later, resulting in a larger spectral distortion with respect to the Fermi-Dirac distribution. Regarding the density distributions for different flavors, we see that electron neutrinos possess larger spectral distortions due to the additional charge-current weak interaction, while the spectral distortions of muon neutrinos almost coincide with those of tauon neutrinos, since they have the same weak interactions. Instead, the spectral distortions of the three neutrino mass eigenstates are uniformly separated, as shown in the right panel.

### v2

@@ 3.1 QKEs in the early universe
In the early universe, both the SM thermal bath and the neutrino system can be approximated as homogeneous and isotropic. Under these conditions, the QKEs for neutrinos can be significantly simplified. First, the drift and force terms on the left-hand side are eliminated, and a Hubble term is introduced to account for the expansion of the Universe. Second, the neutrino density matrix is assumed to depend solely on the absolute value of the momentum. Finally, the coherence matrices between left-handed and right-handed neutrinos, denoted as $\varrho_{LR,RL}$ and $\bar{\varrho}_{LR,RL}$ for Dirac neutrinos, and $\phi$ for Majorana neutrinos, vanish, leading to zero collision terms for these components. Consequently, the non-zero density matrices are $\varrho_{LL,RR}$ and $\bar{\varrho}_{LL,RR}$ for Dirac neutrinos, and $\varrho$ and $\bar{\varrho}$ for Majorana neutrinos.
For active neutrinos, the QKEs for $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ in Dirac neutrinos are identical to those for $\varrho$ and $\bar{\varrho}$ in Majorana neutrinos. Therefore, in the subsequent analysis, we do not distinguish between Dirac and Majorana neutrinos and use $\varrho$ and $\bar{\varrho}$ to represent the density matrices for neutrinos and antineutrinos, respectively. In the flavor basis, the QKEs are written as:
@@
Here, $\mathbb{M}$ represents the neutrino mass matrix in the mass basis, and $U$ is the PMNS matrix, which relates flavor eigenstates to mass eigenstates. Using the standard parameterization of the PMNS matrix Giunti:2007ry ; Gariazzo:2019gyi , we have:
@@
where $c_{ij}=\cos{\theta_{ij}}$, $s_{ij}=\sin{\theta_{ij}}$, and $\theta_{ij}$ are the mixing angles. We neglect the CP-violating phase in this work.
The remaining terms in the Hamiltonian represent matter potentials induced by coherent scattering between neutrinos and other plasma particles. Here, $m_{W}$ and $m_{Z}$ are the masses of the $W$ and $Z$ bosons, respectively, and $G_{F}$ is the Fermi constant. The second term in the Hamiltonian reflects the net number densities of charged leptons ($\mathbb{N}_{l}\equiv\mathrm{diag}(n_{e}-n_{\bar{e}},n_{\mu}-n_{\bar{\mu}},0)$) and neutrinos ($\mathbb{N}_{\nu}\equiv n_{\nu}-n_{\bar{\nu}}$). The third term accounts for the energy densities of charged leptons and neutrinos ($\mathbb{E}_{l}\equiv\mathrm{diag}(\rho_{e}+\rho_{\bar{e}},\rho_{\mu}+\rho_{\bar{\mu}},0)$, $\mathbb{P}_{l}\equiv\mathrm{diag}(P_{e}+P_{\bar{e}},P_{\mu}+P_{\bar{\mu}},0)$, $\mathbb{E}_{\nu}\equiv\rho_{\nu}+\rho_{\bar{\nu}}$). In practice, the net number densities of charged leptons are often negligible due to the small baryon asymmetry. However, the net number densities of neutrinos can significantly affect neutrino flavor oscillations, especially in the presence of primordial neutrino asymmetries. This can result in enhanced neutrino flavor transitions, including the well-known "synchronized oscillations" when this term dominates the Hamiltonian Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa ; Johns:2016enc ; Froustey:2021azz ; Froustey:2024mgf .
Finally, the collision terms $\mathcal{I}$ and $\bar{\mathcal{I}}$ account for contributions from neutrino-electron/positron scattering ($\nu e\leftrightarrow\nu e$), annihilation ($\nu\bar{\nu}\leftrightarrow e^{+}e^{-}$), and neutrino-neutrino self-interactions ($\nu\nu\leftrightarrow\nu\nu$, $\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$). These terms are decomposed as
@@
with a similar decomposition for the antineutrino collision term $\bar{\mathcal{I}}$. The collision terms depend on the neutrino density matrices, and their explicit forms are detailed in Appendix A. Contributions from other processes, such as $\mu^{\pm}$ scattering and annihilation, are neglected due to their minimal impact at the temperatures relevant to neutrino decoupling.
In practice, comoving coordinates and their corresponding variables are employed to eliminate the Hubble expansion term in the QKEs. The comoving temperature is defined as $T_{\rm CM}\propto a^{-1}$ Grohs:2015tfy , with $T_{\rm CM}=T_{\nu}=T_{\gamma}$ at the initial time before neutrino decoupling, when all species are tightly coupled. Using the comoving temperature, the reduced scale factor, comoving momentum, and dimensionless photon temperature are expressed as Mangano:2005cc ; Froustey:2020mcq ; Bennett:2020zkv :
@@
By definition, the dimensionless photon temperature $z$ equals 1 at high temperatures and increases beyond 1 due to the annihilation of electrons and positrons into photons. For convenience in the QKEs, we also define the dimensionless energy density and pressure as $\tilde{\rho}\equiv(x/m_{e})^{4}\rho$ and $\tilde{P}\equiv(x/m_{e})^{4}P$, respectively.
The comoving QKEs for neutrinos and antineutrinos are then written as:
@@
$\frac{{\rm d}\bar{\varrho}(x,y)}{{\rm d}x}$ $=$ $\sqrt{\frac{3m^{2}_{\rm Pl}}{8\pi\tilde{\rho}_{\rm tot}}}\left\{-i\frac{x^{2}}{m_{e}^{3}}\left[\tilde{H}_{\bar{\nu}},\bar{\varrho}\right]+\frac{m_{e}^{3}}{x^{4}}\tilde{\bar{\mathcal{I}}}(\bar{\varrho},\varrho)\right\}\,,$
with the comoving Hamiltonian terms:
@@
Additionally, $\tilde{\mathcal{I}}$ and $\tilde{\bar{\mathcal{I}}}$ represent the comoving collision terms for neutrinos and antineutrinos, respectively. Calculating these collision terms is the most complex and time-consuming aspect of solving the QKEs. Fortunately, the nine-dimensional collision integrals can be reduced to two-dimensional integrals by exploiting the homogeneous and isotropic conditions and the specific forms of the scattering amplitudes Hannestad:1995rs ; Dolgov:1997mb ; Grohs:2015tfy ; Blaschke:2016xxt . The explicit forms of the reduced collision terms are provided in Appendix A.
Furthermore, as noted in the discussion following the QKEs (Eq. (53)), the chemical potential of electrons and the $\mathbb{N}_{l}$ term in the Hamiltonian are negligible due to the very low baryon-to-photon ratio $\eta$ and the efficient electromagnetic interactions. Consequently, the Hamiltonian is predominantly influenced by the net number densities of neutrinos when primordial lepton asymmetries are sufficiently large and vary among different flavors. This results in synchronized flavor oscillations of neutrinos and antineutrinos Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa . As a result, neutrino asymmetries tend to equilibrate among different flavors before the epoch of BBN Froustey:2021azz ; Froustey:2024mgf . Since this work focuses on the implications of neutrino asymmetries for BBN, CMB, and LSS, we concentrate on primordial neutrino asymmetries with complete flavor equilibration before neutrino decoupling. Specifically, the initial primordial asymmetries are set to $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, where $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T_{\rm CM}$ represents the degeneracy parameter for neutrinos and antineutrinos of flavor $\alpha$. By adopting this initial condition, we bypass the intricate mapping between arbitrary primordial neutrino asymmetries and those after flavor equilibration, leaving the investigation of this mapping for future work. Recent studies on this mapping and the implications of distinct neutrino asymmetries among flavors for BBN can be found in Froustey:2024mgf .
On the other hand, it is well known that the annihilation of electrons and positrons shortly after neutrino decoupling partially heats neutrinos, leading to a slightly increased $N_{\rm eff}$. To accurately determine $N_{\rm eff}$, finite-temperature QED (FTQED) corrections must be taken into account Fornengo:1997wa ; Mangano:2001iu ; Bennett:2019ewm . In this work, we address three aspects of FTQED corrections: (1) corrections to the total energy and pressure densities, which primarily affect the Hubble rate; (2) corrections to the electron mass, which mainly influence the collision terms in the QKEs; and (3) corrections to the continuity equation of the Universe, typically expressed as a differential equation for the dimensionless photon temperature $z$,
@@
where $r=x/z$, and $\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}$ and $\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}$ are the diagonal components of the QKEs for neutrinos and antineutrinos. The explicit forms of the remaining $J_{i}(r)$ and $G_{i}(r)$ functions in the equation, as well as other FTQED corrections, can be found in Appendix B. In practice, this equation should be solved in conjunction with the QKEs, and we consider corrections up to order $\mathcal{O}(e^{3})$ while neglecting logarithmic terms in the FTQED corrections, as their effects are minimal Bennett:2019ewm ; Bennett:2020zkv . Further details are provided in Appendix B. Other FTQED corrections, such as those to the interaction vertices, are omitted in this work.
Furthermore, we restrict our analysis to the normal mass hierarchy of neutrinos and neglect any possible CP-violating phase in the transformation matrix between the neutrino flavor and mass bases. Based on the most recent values from the Particle Data Group pdg , we adopt the following neutrino mass differences and mixing angles: $\Delta m_{21}^{2}=7.53\times 10^{-5}\,\text{eV}^{2}$, $\Delta m_{31}^{2}=2.53\times 10^{-3}\,\text{eV}^{2}$, $\sin^{2}\theta_{12}=0.307$, $\sin^{2}\theta_{23}=0.545$, and $\sin^{2}\theta_{13}=0.0218$. The results for neutrino decoupling and BBN are independent of the minimal neutrino mass, provided $m_{1}\ll 0.1\,\text{MeV}$. However, the resulting CMB anisotropies and LSS of the Universe are sensitive to the choice of $m_{1}$, which will be discussed in more detail in Sections 5 and 6.
@@ 3.2 Numerical Methods
To numerically solve the QKEs for neutrinos and antineutrinos, we have extended the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , originally developed for solving QKEs in the Standard Model, to account for primordial neutrino asymmetries. This subsection outlines the primary numerical methods employed in the code.
For the momentum grid, the comoving momentum $y$ is discretized on a linear grid with $y_{\rm min}=0.01$, $y_{\rm max}=20$, and a total of $N_{y}=40$ points. The integrals are evaluated using the first-order Newton-Cotes (NC) formula. The initial conditions are set at an initial reduced scale factor of $x=0.005$, corresponding to a comoving temperature of $T_{\rm CM,in}\simeq 102.2~{}\text{MeV}$. The initial common temperature for all species is calculated by conserving total entropy from a higher temperature, resulting in $z_{\rm in}=2.81\times 10^{-6}$. The initial density matrices are specified as:
$\varrho(x_{\rm in},y)={\rm diag}\{f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y)\},\quad\bar{\varrho}(x_{\rm in},y)={\rm diag}\{f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y)\},$ (60)
@@
$f_{\nu}^{({\rm in})}(y)\equiv\frac{1}{e^{(y-\xi_{\nu})/z_{\rm in}}+1},\quad f_{\bar{\nu}}^{({\rm in})}(y)\equiv\frac{1}{e^{(y+\xi_{\nu})/z_{\rm in}}+1}.$ (61)
The QKEs are solved in conjunction with the energy conservation equation for $z$ using the DLSODA solver from the ODEPACK hindmarsh1982odepack Fortran library. Absolute and relative error tolerances are set to $10^{-7}$, ensuring numerical errors remain below $10^{-5}$.
@@ 3.3 Numerical Results of Neutrino Decoupling
In this subsection, we present the numerical results for neutrino decoupling. First, we focus on the final energy density of neutrinos, typically parameterized by the effective number of relativistic species, $N_{\rm eff}$, defined as
$N_{\rm eff}\equiv\frac{7}{8}\left(\frac{11}{4}\right)^{4/3}\frac{\rho_{\nu}}{\rho_{\gamma}}\,,\quad\rho_{\nu}=\sum_{\alpha}\int\frac{d^{3}p}{(2\pi)^{3}}\,p\,(\varrho_{\nu_{\alpha}}+\bar{\varrho}_{\nu_{\alpha}}),$ (62)
where $\rho_{\nu}$ is the total neutrino energy density, obtained by solving the QKEs, and $\rho_{\gamma}$ is the photon energy density. For the SM in the instantaneous decoupling limit, $N_{\rm eff}$ equals the number of neutrino flavors. However, since the temperature of neutrino decoupling is near that of electron-positron annihilation, neutrinos are also heated through weak interactions with electrons and positrons, leading to a slightly increased $N_{\rm eff}$. The most precise calculation for $N_{\rm eff}$ in the SM, including neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$, and neutrino-electron and neutrino-neutrino collision terms, gives $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
Figure 2: Left panel: Neutrino non-instantaneous decoupling correction $\delta N_{\rm eff}^{\rm dec}=N_{\rm eff}-N_{\rm eff}^{\rm inst}$ as a function of the neutrino degeneracy parameter $\xi_{\nu}$. Solid and dashed curves represent the numerical solution and the fitting function of Eq. (63), respectively. Right panel: Relative differences in the number densities of electron neutrinos and antineutrinos compared to SM predictions in the instantaneous decoupling limit, $n_{\nu_{e},\bar{\nu}_{e}}\simeq 56\,\text{cm}^{-3}$.
In the presence of non-zero primordial neutrino asymmetries, $N_{\rm eff}$ receives additional contributions from the neutrino chemical potentials. In the instantaneous decoupling limit, the comoving distribution of neutrinos and antineutrinos of flavor $\alpha$ is approximated as $f_{\nu_{\alpha},\bar{\nu}_{\alpha}}=\frac{1}{e^{y\,\pm\,\xi_{\nu_{\alpha}}}+1}$. Therefore, the total $N_{\rm eff}$ can be estimated by
$N_{\rm eff}^{\rm inst}\simeq 3+\sum_{\alpha}\left(\frac{30}{7\pi^{2}}\xi_{\nu_{\alpha}}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu_{\alpha}}^{4}\right)=3+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right),$ (63)
@@
We now present the resulting density matrices for neutrinos and antineutrinos after decoupling. The density matrices as a function of the comoving momentum $y\equiv p/T_{\rm CM}$ in both the flavor and mass bases for several values of $\xi_{\nu}$ are shown in the left and right panels of Fig. 3, respectively, and are normalized to the Fermi-Dirac distribution for $\xi_{\nu}=0$. The density matrices for neutrinos and antineutrinos with $\xi_{\nu}=0$ coincide and are represented by the gray curves, reproducing results from previous studies. For non-zero $\xi_{\nu}$, the density matrices of neutrinos and antineutrinos no longer coincide and are represented by red and green curves, respectively. In each case, the solid, dashed, and dash-dotted curves represent the three flavor eigenstates ($\nu_{e}$, $\nu_{\mu}$, and $\nu_{\tau}$) and the three mass eigenstates ($\nu_{1}$, $\nu_{2}$, and $\nu_{3}$), respectively. The corresponding thermal equilibrium distributions for neutrinos and antineutrinos, $f_{\rm FD}^{\nu,\bar{\nu}}=1/(e^{y\pm\xi_{\nu}}+1)$, are also shown by the dotted curves. The density distributions of neutrinos and antineutrinos for a given $\xi_{\nu}$ coincide with the Fermi-Dirac distributions at small $y$, but deviate at larger $y$. This occurs because the weak interaction rate between neutrinos and electrons/positrons is proportional to the neutrino momentum, causing neutrinos with smaller momenta to decouple earlier and maintain the thermal Fermi-Dirac distribution, while neutrinos with larger momenta decouple later, leading to greater spectral distortions relative to the Fermi-Dirac distribution. Regarding different flavors, electron neutrinos exhibit larger spectral distortions due to additional charged-current weak interactions, while the spectral distortions of muon and tau neutrinos nearly coincide, as they experience the same weak interactions. In contrast, the spectral distortions of the three neutrino mass eigenstates are uniformly separated, as shown in the right panel.